OpenStudy (anonymous):

Fan and Medal Solve using perfect square factoring patterns. y^2 + 18y + 81 = 0 A. {9, –9} B. {0, –1} C. {–9} D. {9}

OpenStudy (anonymous):

@atoxicguy

OpenStudy (anonymous):

y^2 + 18y +81 = 0 y^2 + 9y + 9y + 81 = 0 y( y+9) + 9(y+9) = 0 (y+9)*(y+9) = 0 (y+9)^2 = 0 y+9 = 0 and y+9 = 0 => y = -9 in both cases

OpenStudy (anonymous):

isn't that what u r looking?

OpenStudy (anonymous):

so it's C?

OpenStudy (anonymous):

C.